\begin{thebibliography}{23}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\providecommand{\urlprefix}{}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi:\discretionary{}{}{}#1}\else
  \providecommand{\doi}{doi:\discretionary{}{}{}\begingroup
  \urlstyle{rm}\Url}\fi

\bibitem[{Albar and Gon\c{c}alves(2013)}]{BA13}
\textsc{Boris Albar and Daniel Gon\c{c}alves}.
\newblock On triangles in ${K}_r$-minor free graphs.
\newblock 2013.
\newblock \arXiv{1304.5468}.

\bibitem[{Balogh and Kostochka(2011)}]{BK11}
\textsc{J{\'o}zsef Balogh and A.~V. Kostochka}.
\newblock Large minors in graphs with given independence number.
\newblock \emph{Discrete Math.}, 311(20):2203--2215, 2011.
\newblock \doi{10.1016/j.disc.2011.07.003}.

\bibitem[{Balogh et~al.(2011)Balogh, Lenz, and Wu}]{BLW11}
\textsc{J{\'o}zsef Balogh, John Lenz, and Hehui Wu}.
\newblock Complete minors, independent sets, and chordal graphs.
\newblock \emph{Discuss. Math. Graph Theory}, 31(4):639--674, 2011.
\newblock \doi{10.7151/dmgt.1571}.

\bibitem[{Bar{\'a}t et~al.(2011)Bar{\'a}t, Joret, and Wood}]{BJW11}
\textsc{J{\'a}nos Bar{\'a}t, Gwena{\"e}l Joret, and David~R. Wood}.
\newblock Disproof of the list {H}adwiger conjecture.
\newblock \emph{Electron. J. Combin.}, 18(1):P232, 2011.
\newblock
  \urlprefix\url{http://www.combinatorics.org/ojs/index.php/eljc/article/view/v18i1p232}.

\bibitem[{Duchet and Meyniel(1982)}]{DM82}
\textsc{Pierre Duchet and Henri Meyniel}.
\newblock On {H}adwiger's number and the stability number.
\newblock \emph{Annals of Discrete Mathematics}, 13:71--73, 1982.
\newblock \msn{0671905}.

\bibitem[{Fox(2010)}]{Fox10}
\textsc{Jacob Fox}.
\newblock Complete minors and independence number.
\newblock \emph{SIAM J. Discrete Math.}, 24(4):1313--1321, 2010.
\newblock \doi{10.1137/090766814}.

\bibitem[{Hadwiger(1943)}]{Hadwiger43}
\textsc{Hugo Hadwiger}.
\newblock \"{U}ber eine {K}lassifikation der {S}treckenkomplexe.
\newblock \emph{Vierteljschr. Naturforsch. Ges. Z\"urich}, 88:133--142, 1943.
\newblock \MSN{0012237}{6,281c}.

\bibitem[{J{\o}rgensen(1994)}]{Jorgensen94}
\textsc{Leif~K. J{\o}rgensen}.
\newblock Contractions to {$K_8$}.
\newblock \emph{J. Graph Theory}, 18(5):431--448, 1994.
\newblock \doi{10.1002/jgt.3190180502}.

\bibitem[{Kawarabayashi et~al.(2005)Kawarabayashi, Plummer, and Toft}]{KPT05}
\textsc{Ken-ichi Kawarabayashi, Michael~D. Plummer, and Bjarne Toft}.
\newblock Improvements of the theorem of {D}uchet and {M}eyniel on {H}adwiger's
  conjecture.
\newblock \emph{J. Combin. Theory Ser. B}, 95(1):152--167, 2005.
\newblock \doi{10.1016/j.jctb.2005.04.001}.

\bibitem[{Kawarabayashi and Song(2007)}]{KS05}
\textsc{Ken-ichi Kawarabayashi and Zi-Xia Song}.
\newblock Independence number and clique minors.
\newblock \emph{J. Graph Theory}, 56(3):219--226, 2007.
\newblock \doi{10.1002/jgt.20268}.
\newblock \msn{2355127}.

\bibitem[{Kostochka(1982)}]{Kostochka82}
\textsc{Alexandr~V. Kostochka}.
\newblock The minimum {H}adwiger number for graphs with a given mean degree of
  vertices.
\newblock \emph{Metody Diskret. Analiz.}, 38:37--58, 1982.
\newblock \MSN{0713722}{}, \Zbl{0544.05037}.

\bibitem[{Kostochka(1984)}]{Kostochka84}
\textsc{Alexandr~V. Kostochka}.
\newblock Lower bound of the {H}adwiger number of graphs by their average
  degree.
\newblock \emph{Combinatorica}, 4(4):307--316, 1984.
\newblock \doi{10.1007/BF02579141}.
\newblock \MSN{0779891}{86k:05067}.

\bibitem[{Mader(1968)}]{Mader68}
\textsc{Wolfgang Mader}.
\newblock Homomorphies\"atze f\"ur {G}raphen.
\newblock \emph{Math. Ann.}, 178:154--168, 1968.
\newblock \doi{10.1007/BF01350657}.
\newblock \MSN{0229550}{37 \#5124}.

\bibitem[{Maffray and Meyniel(1987)}]{MM-DM87}
\textsc{Fr{\'e}d{\'e}ric Maffray and Henri Meyniel}.
\newblock On a relationship between {H}adwiger and stability numbers.
\newblock \emph{Discrete Math.}, 64(1):39--42, 1987.
\newblock \doi{10.1016/0012-365X(87)90238-X}.
\newblock \msn{882610}.

\bibitem[{Pedersen and Toft(2010)}]{PT10}
\textsc{Anders~Sune Pedersen and Bjarne Toft}.
\newblock A basic elementary extension of the {D}uchet-{M}eyniel theorem.
\newblock \emph{Discrete Math.}, 310(3):480--488, 2010.
\newblock \doi{10.1016/j.disc.2009.03.023}.

\bibitem[{Robertson et~al.(1993)Robertson, Seymour, and Thomas}]{RST-Comb93}
\textsc{Neil Robertson, Paul~D. Seymour, and Robin Thomas}.
\newblock Hadwiger's conjecture for ${K}\sb 6$-free graphs.
\newblock \emph{Combinatorica}, 13(3):279--361, 1993.
\newblock \doi{10.1007/BF01202354}.
\newblock \MSN{1238823}{94i:05037}.

\bibitem[{Song and Thomas(2006)}]{ST06}
\textsc{Zi-Xia Song and Robin Thomas}.
\newblock The extremal function for {$K_9$} minors.
\newblock \emph{J. Combin. Theory Ser. B}, 96(2):240--252, 2006.
\newblock \doi{10.1016/j.jctb.2005.07.008}.
\newblock \MSN{2208353}{2006m:05230}.

\bibitem[{Song(2004)}]{Song04}
\textsc{Zixia Song}.
\newblock \emph{Extremal Functions for Contractions of Graphs}.
\newblock Ph.D. thesis, Georgia Institute of Technology, USA, 2004.
\newblock \cs{10.1.1.92.1275}. \msn{2706190}.

\bibitem[{Thomason(1984)}]{Thomason84}
\textsc{Andrew Thomason}.
\newblock An extremal function for contractions of graphs.
\newblock \emph{Math. Proc. Cambridge Philos. Soc.}, 95(2):261--265, 1984.
\newblock \doi{10.1017/S0305004100061521}.
\newblock \MSN{0735367}{}, \Zbl{0551.05047}.

\bibitem[{Thomason(2001)}]{Thomason01}
\textsc{Andrew Thomason}.
\newblock The extremal function for complete minors.
\newblock \emph{J. Combin. Theory Ser. B}, 81(2):318--338, 2001.
\newblock \doi{10.1006/jctb.2000.2013}.
\newblock \MSN{1814910}{2002i:05100}, \Zbl{1024.05083}.

\bibitem[{Toft(1996)}]{Toft-HadwigerSurvey96}
\textsc{Bjarne Toft}.
\newblock A survey of {H}adwiger's conjecture.
\newblock \emph{Congr. Numer.}, 115:249--283, 1996.
\newblock \MSN{1411244}{97i:05048}.

\bibitem[{Wagner(1937)}]{Wagner37}
\textsc{Klaus Wagner}.
\newblock {\"U}ber eine {E}igenschaft der ebene {K}omplexe.
\newblock \emph{Math. Ann.}, 114:570--590, 1937.
\newblock \doi{10.1007/BF01594196}.
\newblock \msn{1513158}. \Zbl{0017.19005}.

\bibitem[{Wood(2007)}]{Wood-DMTCS07}
\textsc{David~R. Wood}.
\newblock Independent sets in graphs with an excluded clique minor.
\newblock \emph{Discrete Math. Theor. Comput. Sci.}, 9(1):171--175, 2007.
\newblock
  \urlprefix\url{http://www.dmtcs.org/dmtcs-ojs/index.php/dmtcs/article/view/617}.
\newblock \msn{2335892}.

\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{amsmath,amsthm,amsfonts,amssymb,enumerate,calc,iwona}
\usepackage[colorlinks=true,citecolor=black,linkcolor=black,urlcolor=blue]{hyperref}
\usepackage{algorithm}
\usepackage{algorithmic}   
 \floatname{algorithm}{Algorithm}

\usepackage[lmargin=33mm,rmargin=33mm,bmargin=29mm,tmargin=29mm]{geometry}
\usepackage[numbers,sort&compress]{natbib}

\renewcommand{\baselinestretch}{1.1}
\renewcommand{\thefootnote}{\fnsymbol{footnote}}	
\allowdisplaybreaks

\newcommand{\textbold}[1]{{\bf\boldmath #1}}

\newcommand{\Zbl}[1]{Zbl:\,\href{http://www.zentralblatt-math.org/zmath/en/search/?q=an:#1}{#1}}
\newcommand{\arXiv}[1]{arXiv:\,\href{http://arxiv.org/abs/#1}{#1}}
\newcommand{\msn}[1]{MR:\,\href{http://www.ams.org/mathscinet-getitem?mr=MR#1}{#1}}
\newcommand{\MSN}[2]{MR:\,\href{http://www.ams.org/mathscinet-getitem?mr=MR#1}{#1}}
\newcommand{\doi}[1]{doi:\,\href{http://dx.doi.org/#1}{#1}}
\newcommand{\cs}[1]{CiteSeer:\,\href{http://citeseerx.ist.psu.edu/viewdoc/summary?doi=#1}{#1}}

\theoremstyle{plain}
\newtheorem{thm}{Theorem}
\newtheorem{lem}[thm]{Lemma}

\theoremstyle{definition}

\begin{document}

\title{\bf A Note on Hadwiger's Conjecture}

\author{David~R.~Wood\,\footnotemark[1]}

\date{\today}

\maketitle

\footnotetext[1]{School of Mathematical Sciences, Monash University, Melbourne, Australia
  (\texttt{david.wood@monash.edu}). Research supported by  the Australian Research Council.}

\renewcommand{\thefootnote}{\arabic{footnote}}

Hadwiger's Conjecture \citep{Hadwiger43} states that every $K_{t+1}$-minor-free graph is $t$-colourable. It is widely considered to be one of the most important conjectures in graph theory; see \citep{Toft-HadwigerSurvey96} for a survey. If every $K_{t+1}$-minor-free graph has minimum degree at most $\delta$, then every $K_{t+1}$-minor-free graph is $(\delta+1)$-colourable by a minimum-degree-greedy algorithm. The purpose of this note is to prove a slightly better upper bound.

\begin{lem}
\label{Main}
Fix $t\geq 2$. Assume that every $K_{t+1}$-minor-free graph has minimum degree at most $\delta$, and that every $K_{t}$-minor-free  graph with exactly $\delta$ vertices has an independent set of size $\alpha$. Then every $K_{t+1}$-minor-free graph is $(\delta-\alpha+2)$-colourable.
\end{lem}

\begin{proof}
We prove, by induction on $n$, that every $n$-vertex $K_{t+1}$-minor-free graph is $(\delta-\alpha+2)$-colourable. 
The base case with $n=1$ is trivial since $\delta-\alpha+2\geq 2$. 
Let $G$ be a $K_{t+1}$-minor-free graph. 
Let $v$ be a vertex of minimum degree $d$ in $G$. 
Thus $d\leq \delta$. 
If $d=0$ then, by induction, $G-v$ is $(\delta-\alpha+2)$-colourable, and by assigning to $v$ any colour already in use, we obtain a $(\delta-\alpha+2)$-colouring of $G$, as desired. 
Now assume that $d\geq 1$. 

Let $H$ be the subgraph of $G$ induced by $N(v)$. 
Thus $H$ has $d$ vertices and no $K_t$-minor. 
Let $H'$ be the graph obtained from $H$ by adding $\delta-d$ isolated vertices. 
Then $H'$ has exactly $\delta$ vertices and $H'$ also has no $K_t$-minor. 
By assumption, $H'$ has an independent set of size $\alpha$. 
Thus $H$ has an independent set $T$ of size $\alpha-\delta+d$. 

Let $G'$ be the graph obtained from $G$ by contracting each edge $vw$ where $w\in T$ into a new vertex $z$. 
Since $G'$ is a minor of $G$, $G'$ is $K_{t+1}$-minor-free. Since $d\geq 1$, $G'$ has less vertices than $G$. 
By induction, $G'$ is $(\delta-\alpha+2)$-colourable. 
Colour each vertex in $T$ by the colour assigned to $z$.
Colour each vertex in $G-T-v$ by the colour assigned to the same vertex in $G'$. 
Of the $d$ neighbours of $v$, at least $\alpha-\delta+d$ have the same colour. 
Thus at most $d-(\alpha-\delta+d)+1=\delta-\alpha+1$ colours are present on the neighbours of $v$. 
Hence, at least one of the $\delta-\alpha+2$ colours is not assigned to a neighbour of $v$, and this colour may be assigned to $v$. Thus $G$ is $(\delta-\alpha+2)$-colourable.
\end{proof}

The next lemma summarises some results about independent sets in a $K_{t+1}$-minor-free graphs.
Part (a) is the original result in this direction by \citet{DM82}.
Part (b), which is strong when $t$ is small, is by the author \citep{Wood-DMTCS07}. 
Part (c), which builds upon a similar result by \citet{Fox10}, is due to \citet{BK11}. See  \citep{MM-DM87,BLW11,PT10,KS05,KPT05} for related results. 

\begin{lem}
\label{BigIndSet}
Every $K_{t+1}$-minor-free graph on $n$ vertices has an independent set of size $\alpha$, where  \\
\hspace*{5mm} \textup{(a)} \quad  $(2\alpha -1)t \geq n$\enspace,\\
\hspace*{5mm} \textup{(b)} \quad  $(2\alpha -1)(2t- 5) \geq 2n -5$ for $t\geq 5$\enspace,\\
\hspace*{5mm} \textup{(c)} \quad  $(2-\gamma)\alpha t\geq n$\enspace, \\
where $\gamma= (80-\sqrt{5392})/126 \approx  0.0521\dots$.
\end{lem}

\begin{thm}
\label{Corollary}
Fix $t\geq 6$. Assume that every $K_{t+1}$-minor-free graph $G$ has minimum degree at most $\delta$.
Then $$\chi(G)\leq \delta- \frac{2\delta -5}{4t - 14}+\frac{3}{2}$$
and $$\chi(G)\leq\delta\left(1-\frac{1}{(2-\gamma)(t-1)}\right)+2\enspace,$$
where $\gamma= (80-\sqrt{5392})/126 \approx  0.0521\dots$.
\end{thm}

\begin{proof}
Lemma~\ref{BigIndSet}(b) implies that every $K_t$-minor-free graph with $\delta$ vertices has an independent set of size $\alpha$, where $(2\alpha -1)(2(t-1)- 5) \geq 2\delta -5$. Thus $\alpha\geq \frac{2\delta-5}{4t-14}+\frac{1}{2}$. Lemma~\ref{Main} implies that $\chi(G)\leq \delta-\alpha+2\leq \delta-\frac{2\delta-5}{4t-14}+\frac{3}{2}$. Similarly, Lemma~\ref{BigIndSet}(c) implies the second result. 
\end{proof}

Note that \citet{Kostochka82,Kostochka84} and \citet{Thomason01,Thomason84} independently proved that  $\delta\leq ct\sqrt{\log t}$ is the best possible upper bound on the minimum degree of $K_{t+1}$-minor-free graphs. Thus such graphs are $ct\sqrt{\log t}$-colourable. Unfortunately, Theorem~\ref{Corollary} makes no asymptotic improvement to this result. 

\medskip
We now apply these results for particular values of $t$. 

\textbold{$t=2$:} $K_3$-minor-free graphs are exactly the forests, and every forest has a vertex of degree 1. Thus $\chi(G)\leq 1-1+2=2$  by Lemma~\ref{Main}, which is tight. 

\textbold{$t=3$:}  Every $K_4$-minor-free graph $G$ has minimum degree at most 2, and every 2-vertex graph has an independent set of size 1. Thus $\chi(G)\leq 2-1+2=3$  by Lemma~\ref{Main}, which is tight. 

\textbold{$t=4$:}  Every $K_5$-minor-free graph with at least 3 vertices has at most $3n-6$ edges \citep{Mader68}. 
Thus every $K_5$-minor-free graph  has average degree less than 6 and minimum degree at most 5. Every 5-vertex $K_4$-minor-free graph has an independent set of size 2. Thus $\chi(G)\leq 5-2+2=5$  by Lemma~\ref{Main}. The 4-colour theorem and Wagner's characterisation \cite{Wagner37} implies that $\chi(G)\leq 4$. 

\textbold{$t=5$:}  Every $K_6$-minor-free graph with at least 4 vertices has at most $4n-10$ edges \citep{Mader68}. Thus Every $K_6$-minor-free graph has average degree less than 8, and minimum degree at most 7. Every 7-vertex $K_5$-free graph has an independent set of size $2$. Thus every $K_6$-minor-free graph is 7-colourable by Lemma~\ref{Main}, which is inferior to the result by \citet{RST-Comb93} who proved that such graphs are 5-colourable. Note that it is open whether  every $K_6$-minor-free graph has minimum degree at most 6 (see \citep{BJW11}). 

\textbold{$t=6$:}  Every $K_7$-minor-free graph with at least $5$ vertices has at most $5n-15$ edges \citep{Mader68}.
Thus every $K_7$-minor-free graph  has average degree less than 10, and minimum degree at most 9. Every 9-vertex $K_6$-free graph has an independent set of size $2$. Thus every $K_7$-minor-free graph is 9-colourable  by Lemma~\ref{Main}. 
\citet{BA13} proved that every $K_7$-minor-free graph is 8-colourable. 

We conjecture that every $K_7$-minor-free graph has minimum degree at most $7$ (which would be tight for $K_{1,2,2,2,2}$). This conjecture would imply that every $K_7$-minor-free graph is 7-colourable. Hadwiger's conjecture says that $K_7$-minor-free graphs are 6-colourable. 

\textbold{$t=7$:}  Every $K_8$-minor-free graph with at least 5 vertices has at most $6n-20$ edges \citep{Jorgensen94}, 
Thus every $K_8$-minor-free graph has average degree less than 12, and minimum degree at most 11. Every 11-vertex $K_7$-free graph has an independent set of size $2$. Thus every $K_8$-minor-free graph is 11-colourable by Lemma~\ref{Main}. \citet{BA13} proved that every $K_7$-minor-free graph is $10$-colourable.

\citet{Jorgensen94} characterised the $K_8$-minor-free graphs with $6n-20$ edges as those obtained from copies of $K_{2,2,2,2,2}$ by pasting on 5-cliques. Such graphs have minimum degree 8. We conjecture that every $K_8$-minor-free has minimum degree at most 8, which would imply that such graphs are 8-colourable. Hadwiger's conjecture says that $K_8$-minor-free graphs are 7-colourable. 

\textbold{$t=8$:}  Every $K_9$-minor-free graph $G$ has at most $7n-27$ edges \citep{ST06}, has average degree less than 14, and minimum degree at most 13. Every 13-vertex $K_8$-free graph has an independent set of size $2$. Thus 
every $K_9$-minor-free graph is $13$-colourable by Lemma~\ref{Main}. 

\citet{ST06} characterised the $K_9$-minor-free graphs with $7n-27$ edges as those obtained from copies of $K_{2,2,2,3,3}$ and $K_{1,2,2,2,2,2}$ by pasting on 6-cliques. Such graphs have minimum degree at most 10. We conjecture that every $K_9$-minor-free has minimum degree at most 10, which would imply that such graphs are 10-colourable by Lemma~\ref{Main}. Hadwiger's conjecture says that $K_9$-minor-free graphs are 8-colourable. 

\textbold{$t=9$}:  Every $K_{10}$-minor-free graph with at least 10 vertices has at most $11n- 66$ edges \citep{Song04}. Thus every $K_{10}$-minor-free graph has average degree less than $22$, and minimum degree at most 21. By Lemma~\ref{BigIndSet}(b), every $K_9$-minor-free 21-vertex graph has an independent set of 3 vertices. Thus  every $K_{10}$-minor-free graph is $20$-colourable by Lemma~\ref{Main} (or directly by Theorem~\ref{Corollary}). 

\textbold{$t=10$:}  Every $K_{11}$-minor-free graph with at least 11 vertices has at most $13n- 89$ edges \citep{Song04}. Thus every $K_{11}$-minor-free graph has average degree less than $26$, and minimum degree at most 25. By Lemma~\ref{BigIndSet}(b), every $K_{10}$-minor-free 25-vertex graph has an independent set of 3 vertices. Thus  every $K_{11}$-minor-free graph is $24$-colourable by Lemma~\ref{Main} (or directly by Theorem~\ref{Corollary}). 

We emphasis that in the cases $t=9$ and $t=10$, while it is likely that all of the stated bounds are far from optimal\footnote{See Chapter 6 of the Ph.D.\ thesis of \citet{Song04} for a discussion of the likely extremal graphs}, the utility of our approach is evident, since we may take $\alpha=3$ in these cases. 


\def\soft#1{\leavevmode\setbox0=\hbox{h}\dimen7=\ht0\advance \dimen7
  by-1ex\relax\if t#1\relax\rlap{\raise.6\dimen7
  \hbox{\kern.3ex\char'47}}#1\relax\else\if T#1\relax
  \rlap{\raise.5\dimen7\hbox{\kern1.3ex\char'47}}#1\relax \else\if
  d#1\relax\rlap{\raise.5\dimen7\hbox{\kern.9ex \char'47}}#1\relax\else\if
  D#1\relax\rlap{\raise.5\dimen7 \hbox{\kern1.4ex\char'47}}#1\relax\else\if
  l#1\relax \rlap{\raise.5\dimen7\hbox{\kern.4ex\char'47}}#1\relax \else\if
  L#1\relax\rlap{\raise.5\dimen7\hbox{\kern.7ex
  \char'47}}#1\relax\else\message{accent \string\soft \space #1 not
  defined!}#1\relax\fi\fi\fi\fi\fi\fi}

\end{document}